# Harmonizing SO(3)-Equivariance with Neural Expressiveness: a Hybrid Deep Learning Framework Oriented to the Prediction of Electronic Structure Hamiltonian

### v9

@@ 5.2 Experimental Comparison and Analysis
In addition to taking the Hamiltonian of each edge as a whole for accuracy statistics, since the Hamiltonian matrix in the direct product state is constituted by several basic blocks based on the angular momentum of interacting orbitals, we also conduct fine-grained accuracy statistics on these basic blocks. The MAE metrics (denoted as $MAE_{block}$) of our method and DeepHE3 on different blocks of the Hamiltonian matrix of the six structures are presented in Fig. 3 and 4, where Fig. 3 illustrates the results on monolayer structures, while Fig. 4 is dedicated to the bilayer structures. These results of $MAE_{block}$ in these tables are gathered from the mean of $10$ independent experiments.
Table 2: Comparison of experimental results on the monolayer structures. Both the mean and standard deviation of $MAE_{all}$ as well as $MAE_{cha}$ (/meV) from $10$ independent repeated experiments are provided.
Dataset Method $MAE_{all}$ $MAE_{cha}$
MG $DeepH\!E3$ 0.251 $\pm$ 0.002 0.357 $\pm$ 0.002
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.239 $\pm$ 0.003 0.338 $\pm$ 0.003
$GFormer$ 0.816 $\pm$ 0.005 0.897 $\pm$ 0.005
$Stage1_{GFormer}+Stage2_{GFormer}$ 0.653 $\pm$ 0.004 0.720 $\pm$ 0.005
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 0.774 $\pm$ 0.005 0.880 $\pm$ 0.006
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.243 $\pm$ 0.001 0.328 $\pm$ 0.002
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.221 $\pm$ 0.002 0.297 $\pm$ 0.001
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.176 $\pm$ 0.001 0.267 $\pm$ 0.001
MM $DeepH\!E3$ 0.406 $\pm$ 0.002 0.574 $\pm$ 0.003
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.392 $\pm$ 0.001 0.499 $\pm$ 0.002
$GFormer$ 1.025 $\pm$ 0.004 1.250 $\pm$ 0.004
$Stage1_{GFormer}+Stage2_{GFormer}$ 0.911 $\pm$ 0.003 0.923 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 0.927 $\pm$ 0.004 0.958 $\pm$ 0.003
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.384 $\pm$ 0.002 0.414 $\pm$ 0.002
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.319 $\pm$ 0.001 0.366 $\pm$ 0.002
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.233 $\pm$ 0.001 0.293 $\pm$ 0.002
Dataset Method $MAE_{all}$ $MAE_{cha}$
BGnt $DeepH\!E3$ 0.389 $\pm$ 0.002 0.453 $\pm$ 0.003
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.372 $\pm$ 0.001 0.434 $\pm$ 0.002
$GFormer$ 1.295 $\pm$ 0.005 1.483 $\pm$ 0.004
$Stage1_{GFormer}+Stage2_{GFormer}$ 0.786 $\pm$ 0.005 0.828 $\pm$ 0.005
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 0.854 $\pm$ 0.003 0.920 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.365 $\pm$ 0.002 0.427 $\pm$ 0.002
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.348 $\pm$ 0.001 0.419 $\pm$ 0.003
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.287 $\pm$ 0.002 0.362 $\pm$ 0.002
BBnt $DeepH\!E3$ 0.274 $\pm$ 0.003 0.304 $\pm$ 0.002
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.268 $\pm$ 0.004 0.292 $\pm$ 0.003
$GFormer$ 0.886 $\pm$ 0.004 0.949 $\pm$ 0.005
$Stage1_{GFormer}+Stage2_{GFormer}$ 0.785 $\pm$ 0.004 0.816 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 0.802 $\pm$ 0.003 0.873 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.249 $\pm$ 0.004 0.281 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.243 $\pm$ 0.001 0.286 $\pm$ 0.003
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.172 $\pm$ 0.001 0.198 $\pm$ 0.002
BTnt $DeepH\!E3$ 0.447 $\pm$ 0.001 0.480 $\pm$ 0.002
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.435 $\pm$ 0.002 0.471 $\pm$ 0.002
$GFormer$ 1.018 $\pm$ 0.005 1.230 $\pm$ 0.006
$Stage1_{GFormer}+Stage2_{GFormer}$ 0.903 $\pm$ 0.003 0.982 $\pm$ 0.005
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 0.930 $\pm$ 0.004 1.026 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.439 $\pm$ 0.001 0.466 $\pm$ 0.003
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.385 $\pm$ 0.001 0.414 $\pm$ 0.002
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.294 $\pm$ 0.002 0.321 $\pm$ 0.002
BSnt $DeepH\!E3$ 0.397 $\pm$0.003 0.424 $\pm$0.001
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.389 $\pm$ 0.002 0.410 $\pm$ 0.002
$GFormer$ 1.352 $\pm$ 0.005 1.483 $\pm$ 0.005
$Stage1_{GFormer}+Stage2_{GFormer}$ 0.862 $\pm$ 0.005 0.922 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 0.898 $\pm$ 0.005 0.960 $\pm$ 0.005
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.392 $\pm$ 0.002 0.401 $\pm$ 0.001
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.348 $\pm$ 0.002 0.375 $\pm$ 0.003
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.282 $\pm$ 0.002 0.308 $\pm$ 0.002
Dataset Method $MAE_{all}$ $MAE_{cha}$
BGt $DeepH\!E3$ 0.264 $\pm$ 0.003 0.429 $\pm$ 0.003
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.257 $\pm$ 0.002 0.423 $\pm$ 0.003
$GFormer$ 0.982 $\pm$ 0.005 1.153 $\pm$ 0.006
$Stage1_{GFormer}+Stage2_{GFormer}$ 0.841 $\pm$ 0.004 0.873 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 0.801 $\pm$ 0.005 0.863 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.312 $\pm$ 0.003 0.441 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.278 $\pm$ 0.002 0.428 $\pm$ 0.002
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.227 $\pm$ 0.002 0.403 $\pm$ 0.001
BBt $DeepH\!E3$ 0.468 $\pm$ 0.003 0.602 $\pm$0.004
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.460 $\pm$ 0.003 0.595 $\pm$0.004
$GFormer$ 1.784 $\pm$ 0.006 1.921 $\pm$ 0.006
$Stage1_{GFormer}+Stage2_{GFormer}$ 1.426 $\pm$ 0.005 1.680 $\pm$ 0.006
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 1.213 $\pm$ 0.004 1.569 $\pm$ 0.005
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.530 $\pm$ 0.004 0.697 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.504 $\pm$ 0.004 0.669 $\pm$ 0.005
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.438 $\pm$0.003 0.578 $\pm$0.002
BTt $DeepH\!E3$ 0.831 $\pm$ 0.003 0.850 $\pm$ 0.003
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.826 $\pm$ 0.002 0.843 $\pm$ 0.002
$GFormer$ 2.682 $\pm$ 0.004 2.827 $\pm$ 0.006
$Stage1_{GFormer}+Stage2_{GFormer}$ 2.190 $\pm$ 0.003 2.379 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 1.892 $\pm$ 0.003 1.937 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.928 $\pm$ 0.004 0.943 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.837 $\pm$ 0.003 0.846 $\pm$ 0.003
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.774 $\pm$ 0.002 0.794 $\pm$ 0.003
BSt $DeepH\!E3$ 0.370 $\pm$ 0.001 0.390 $\pm$ 0.002
$Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ 0.358 $\pm$ 0.001 0.381 $\pm$ 0.002
$GFormer$ 1.785 $\pm$ 0.004 2.037 $\pm$ 0.005
$Stage1_{GFormer}+Stage2_{GFormer}$ 1.624 $\pm$ 0.004 1.953 $\pm$ 0.004
$Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$ 1.569 $\pm$ 0.003 1.892 $\pm$ 0.005
$Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$ 0.415 $\pm$ 0.003 0.456 $\pm$ 0.003
$Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$ 0.392 $\pm$ 0.002 0.421 $\pm$ 0.002
$Ours@(Stage1_{DeepH\!E3}+Stage2_{GFormer})$ 0.336 $\pm$ 0.001 0.365 $\pm$ 0.002
From these tables and figures, we find that the proposed hybrid approach significantly improves the prediction accuracy beyond what is achievable solely with the parameter-independently SO(3)-equivariant mechanisms in $DeepH\!E3$ or the non-linear mechanisms in $GFormer$, demonstrating the effectiveness on leveraging the complementarity of the two categories of neural mechanisms to overcome their respective challenges comprehensively analyzed in previous sections. On one hand, the highly expressive non-linear mechanisms in the Transformer effectively compensate for the limitations in non-linear expressiveness of $DeepH\!E3$’s mechanisms, thus lowering down the $MAE_{all}$ of $DeepH\!E3$ by a relative ratio of $29.88\%$, $42.61\%$, $26.22\%$, $37.23\%$, $34.23\%$, $28.97\%$, $14.02\%$, $6.41\%$, $6.86\%$, and $9.19\%$, on MG, MM, BGnt, BBnt, BTnt, BSnt, BGt, BBt, BTt and BSt, respectively; and lowering down $MAE_{cha}$ by $25.21\%$, $48.95\%$, $20.09\%$, $34.87\%$, $33.13\%$, $27.36\%$, $6.06\%$, $3.99\%$, $6.59\%$, and $6.41\%$, on the respective datasets; and achieving better results in the vast majority of basic blocks, especially for the blocks where $DeepH\!E3$ performs the worst, by reducing $MAE_{block}$ of these challenging blocks by $13.89\%$, $66.36\%$, $29.69\%$, $39.42\%$, $33.81\%$, $28.74\%$, $13.11\%$, $10.83\%$, $5.03\%$ and $7.95\%$, on the respective datasets, from which the $MAE_{block}$ of challenging blocks in MM, BGnt and BTnt are surprisingly lowering down to within $1mev$. On the other hand, the mechanisms of $DeepH\!E3$ help the non-linear mechanisms in the Transformer to better learn equivariance from the data and reduces the difficulty for the Transformer on regressing SO(3)-equivariant targets, thus also significantly promoting the results from merely using the Transformer network. From the experimental results, it is observed that without any prior information on covariance, solely relying on the non-linear Transformer architecture to learn SO(3)-equivariance from data is extremely challenging despite rotational augmentation, due to the complexity and high-dimensionality nature of Hamiltonians as shown in Table 1. And since SO(3)-equivariance is strongly linked to the intrinsic mathematical structure of Hamiltonians, the weakness on capturing SO(3)-equivariance results in inadequate modeling of Hamiltonians, leading to inaccurate predictions, especially for samples that exhibit obvious SO(3)-equivariant effects, such as twisted samples. Our framework addresses this challenge by incorporating mechanisms with prior covariance, which helps the non-linear mechanisms in the Transformer to learn equivariance from the data and reduces the difficulty for non-linear regression of SO(3)-equivariant targets, bringing satisfactory performance of the non-linear modules. The results on twisted samples demonstrate that, our method, despite the extensive use of non-linear operators, can still capture the symmetry properties of Hamiltonians and make equivariant predictions under rotational operations. In contrast to this, the experimental results from $Stage1_{DeepH\!E3}+Stage2_{DeepH\!E3}$ and $Stage1_{GFormer}+Stage2_{GFormer}$ shows that simply scaling up the parameters for $DeepH\!E3$ or $GFormer$ and fine-tuning it alone only yields limited improvements. This indicates that the bottlenecks encountered by these two categories of neural mechanisms might not be fully overcome through scaling up their sizes, further highlighting the superiority and necessity of our hybrid strategy.
Figure 3: The MAE values (denoted as $MAE_{\text{block}}$) on various basic blocks of the Hamiltonian matrix in direct product state on the experimental monolayer structures.
Figure 4: The MAE values (denoted as $MAE_{\text{block}}$) on various basic blocks of the Hamiltonian matrix in direct product state on the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) bilayer structures.
As fine-grained ablation studies, by comparing the results of the three experimental settings, i.e., $Stage1_{DeepH\!E3}+Stage2^{-cas}_{GFormer}$, $Stage1_{DeepH\!E3}+Stage2^{-cov}_{GFormer}$, and $Stage1_{DeepH\!E3}+Stage2^{-att}_{GFormer}$, with our complete method, we could observe that the cascaded regression mechanism, the covariant feature integration mechanism, as well the multi-head attention mechanism, all contribute significantly to the performance of our method. The cascaded regression mechanism, by reducing the output space of the non-linear network, eases the difficulties on non-linear regression of Hamiltonians with SO(3)-equivariance; the covariant feature integration mechanism, through leveraging theoretical-guaranteed covariant features from DeepHE3 and geometric knowledge, successfully assists the non-linear network in learning SO(3)-equivariance; the multi-head attention mechanism, by assigning dynamic weights when fusing features, adapts to the wide variation range of geometric conditions, including both thermal deformations and twists. Under the novel combination of these mechanisms, networks from the two stage complement each other effectively, making our framework possess both excellent expressive capability and SO(3)-equivariant performance to achieve good results.
Among current deep learning studies on Hamiltonians modeling, the DeepH series have conducted comprehensive experiments on various types of atomic structures, proving the SOTA performance. The paper of DeepHE3 has demonstrated it to surpass DeepH in accuracy by integrating neural mechanisms with prior equivariance; and since our method effectively harmonizes SO(3)-equivariance and expressiveness of deep neural networks, we further surpass DeepHE3 in all of the benchmark databases, both on average and for the most challenging samples and blocks, both on non-twisted and twisted samples, conforming the leading performance in the field of deep Hamiltonian regression.
6 Time Complexity
The time complexity for both the first-stage and the second-stage networks is $\mathcal{O}(BN\overline{E})$, where $B$ is the number of their basic blocks, $N$ is the total number of atoms in a system, and $\overline{E}$ is the average number of neighboring atoms within the cutoff radius of a given atom. When $N$ is small, the system forms edges between each pair of atoms, and $\overline{E}$ is approximately equal to $N$. However, as $N$ increases to a large value, due to the locality of the Hamiltonian, the set of neighboring atoms for an atom typically remains roughly stable, so $\overline{E}$ is roughly a constant and does not significantly increase with the growth of $N$. Therefore, for big atomic systems with very large $N$, $\overline{E}\ll N$, and similarly, $B\ll N$. In this case, the time complexity of our framework tends towards $\mathcal{O}(N)$. In contrast, for the traditional DFT method [11], which requires $T$ rounds of iterative decomposition of $N\times N$ matrices, the time complexity is $\mathcal{O}(TN^{3})$. As $N$ increases, the cubic complexity of $N$ results in significant computational overhead, and to achieve self-consistency, $T$ is usually very large, exacerbating the computational burden, leading to difficulties for simulations of large atomic systems within an acceptable time frame. On the contrary, our framework offers an efficient pathway for simulating large atomic systems with linear time complexity of $N$ and without the need for $T$ rounds of iterations.
7 Conclusion
Deep learning for regressing electronic structure Hamiltonian requires adherence to covariance laws, facing a pivotal challenge to achieve SO(3)-equivariance without compromising neural expressiveness. To solve this, we make an exploration towards highly-expressive SO(3)-equivariant deep learning by proposing a hybrid two-stage regression framework, where the first stage employs neural mechanisms inherent with SO(3)-equivariance properties prior to the learning process based on group theory, yielding baseline Hamiltonians with series of equivariant features assisting the subsequent stage on capturing SO(3)-equivariance. The second stage, leveraging the proposed non-linear 3D graph Transformer network for fine-grained structural analysis of 3D atomic systems, learns SO(3)-equivariant patterns from training data with the help of the first stage, while in turn, refines the initial Hamiltonian predictions via enhanced network expressiveness. Such a combination allows for accurate, generalizable Hamiltonian predictions while upholding strong SO(3)-equivariance against rotational transformations. Our methodology demonstrates SOTA performance in Hamiltonian prediction, validated through six benchmark databases, showing good potentials in high-performance simulation of atomic systems.
8 Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant Number 12134012) and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant Number XDB0500201).

### v11

@@ Introduction
Experimental results of our complete method as well as the compared experimental settings on the six benchmark databases are presented in Tables 2, 3, and 4, respectively detailing the results for monolayer structures, as well as the the results for non-twisted and twisted samples of bilayer structures. In these tables, the Mean Absolute Error (MAE) metric is used as the accuracy metric. Besides the classical MAE metric, denoted as $MAE^{H}_{all}$, which measures the average error among all testing samples, we also record $MAE^{H}_{cha\_s}$, the MAE for the most challenging sample where the baseline (DeepHE3) performs the worst. In addition to taking the Hamiltonian of each edge as a whole for accuracy statistics, since the Hamiltonian matrix in the direct product state is constituted by several basic blocks based on the angular momentum of interacting orbitals, we also conduct fine-grained accuracy statistics on these basic blocks. The MAE metrics (denoted as $MAE^{H}_{block}$) of our method and DeepHE3 on different blocks of the Hamiltonian matrix of the six structures are presented in Fig. 3 and 4, where Fig. 3 illustrates the results on monolayer structures, while Fig. 4 is dedicated to the bilayer structures. In these Figures, we specifically highlight the MAE values (denoted as $MAE^{H}_{cha\_b}$) on the Hamiltonian block where the baseline model DeepHE3 performs the worst for comparison on challenging blocks. All presented results are the mean values from $10$ independent repeat experiments. Since a fixed random seed is used, the standard deviation is less than 0.008 meV and negligible.
Figure 4: The MAE values (denoted as $MAE^{H}_{block}$) on various basic blocks of the Hamiltonian matrix in direct product state on the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) bilayer structures. The MAE values ($MAE^{H}_{cha\_b}$) on the Hamiltonian block where the baseline model DeepHE3 performs the worst are highlighted for comparison.
Observed from Table 2, 3, 4, and Figures 1, 2, we find that the proposed hybrid approach significantly improves the prediction accuracy beyond what is achievable solely with the priorly-equivariant mechanisms in $DeepH\!E3$ or the non-linear mechanisms in $GFormer$, demonstrating the effectiveness on leveraging the complementarity of the two categories of neural mechanisms to overcome their respective challenges comprehensively analyzed in previous sections. On one hand, the highly expressive non-linear mechanisms in the Transformer effectively compensate for the limitations in non-linear expressiveness of $DeepH\!E3$’s mechanisms, thus significantly lowering down the $MAE^{H}_{all}$ and $MAE^{H}_{cha_{s}}$ of $DeepHE3$, and decreasing $MAE^{H}_{block}$ for the vast majority of basic blocks, particularly for those blocks where $DeepHE3$ performs the worst. On the other hand, the mechanisms of $DeepH\!E3$ help the non-linear mechanisms in the Transformer to better learn equivariance from the data and reduces the difficulty for the Transformer on regressing SO(3)-equivariant targets, thus also significantly promoting the results from merely using the Transformer network. From the experimental results, it is observed that without any prior information on covariance, solely relying on the non-linear Transformer architecture to learn SO(3)-equivariance from data is extremely challenging despite rotational augmentation, due to the complexity and high-dimensionality nature of Hamiltonians as shown in Table 1. And since SO(3)-equivariance is strongly linked to the intrinsic mathematical structure of Hamiltonians, the weakness on capturing SO(3)-equivariance results in inadequate modeling of Hamiltonians, leading to inaccurate predictions, especially for samples that exhibit obvious SO(3)-equivariant effects, such as twisted samples. Our framework mitigates this challenge by incorporating mechanisms with prior covariance, which helps the non-linear mechanisms in the Transformer to learn equivariance from the data and reduces the difficulty for non-linear regression of SO(3)-equivariant targets, bringing satisfactory performance of the non-linear modules. The results on twisted samples demonstrate that, our method, despite the extensive use of non-linear operators, can still capture the symmetry properties of Hamiltonians and make equivariant predictions under rotational operations. In contrast to this, the experimental results from $S1_{DeepH\!E3}+S2_{DeepH\!E3}$ and $S1_{GFormer}+S2_{GFormer}$ shows that simply scaling up the parameters for $DeepH\!E3$ or $GFormer$ and fine-tuning it alone only yields limited improvements. This indicates that the bottlenecks encountered by these two categories of neural mechanisms might not be fully overcome through scaling up their sizes, further highlighting the superiority and necessity of our hybrid framework.
As fine-grained ablation studies, by comparing the results of the three experimental settings, i.e., $S1_{DeepH\!E3}+S2^{-cas}_{GFormer}$, $S1_{DeepH\!E3}+S2^{-cov}_{GFormer}$, and $S1_{DeepH\!E3}+S2^{-att}_{GFormer}$, with our complete method, we could observe that the cascaded regression mechanism, the covariant feature integration mechanism, as well the multi-head attention mechanism, all contribute significantly to the performance of our method. The cascaded regression mechanism, by reducing the output space of the non-linear network, eases the difficulties on non-linear regression of Hamiltonians with SO(3)-equivariance; the covariant feature integration mechanism, through leveraging theoretical-guaranteed covariant features from DeepHE3 and geometric knowledge, successfully assists the non-linear network in learning SO(3)-equivariance; the multi-head attention mechanism, by assigning dynamic weights when fusing features, adapts to the wide variation range of geometric conditions, including both thermal deformations and twists. Under the combination of these mechanisms, networks from the two stage complement each other effectively, making our framework possess both excellent expressive capability and SO(3)-equivariant performance to achieve good results.
Conclusion
Deep learning for regressing electronic-structure Hamiltonian faces a pivotal challenge to capture SO(3)-equivariance without compromising neural expressiveness. To solve this, we propose a hybrid two-stage encoding and regression framework, where the first stage employs neural mechanisms inherent with SO(3)-equivariance properties prior to the learning process based on group theory, yielding baseline Hamiltonians with series of equivariant features assisting the subsequent stage on capturing SO(3)-equivariance. The second stage, leveraging the proposed non-linear 3D graph Transformer network for fine-grained structural analysis of 3D atomic systems, learns SO(3)-equivariant patterns from training data with the help of the first stage, while in turn, refines the initial Hamiltonian predictions via enhanced network expressiveness. Such a combination allows for accurate, generalizable Hamiltonian predictions while upholding good equivariant performance against rotational transformations. Our methodology demonstrates SOTA performance in Hamiltonian prediction, validated through six benchmark databases, showing good potentials in high-performance deep modeling of atomic systems.